# Iterative Prompt Relocation for Distribution-Adaptive Visual Prompt Tuning

### v1

@@ Appendix A Detailed Explanations of Technical Concepts
where $\pi$ is the policy, $\theta$ is the policy parameter, and $t$ is the $t^{\text{th}}$ step. It typically takes multiple steps of SGD to optimize the objective function $L$:
$L(s,a,\theta_{t},\theta)=$
$\;\min\big{(}\frac{\pi_{\theta}(a\!\mid\!s)}{\pi_{\theta_{t}}(a\!\mid\!s)}A^{\pi\theta_{t}}(s,a),\text{clip}(\frac{\pi_{\theta}(a\!\mid\!s)}{\pi_{\theta_{t}}(a\!\mid\!s)},1\!-\!\epsilon,1\!+\!\epsilon)A^{\pi\theta_{t}}(s,a)\big{)},$ (vii)
where $A^{\pi\theta_{t}}(s,a)$ is the advantage estimator and $\epsilon$ is the clip hyper-parameter. The clip term restricts the policy change, effectively preventing the new policy from deviating significantly from the old policy [33].
Appendix B Comprehensive Analysis of the Underlying Nature behind ADO
@@
Fig. i presents the complete results from Fig. 2. Specifically, we investigated the performance gaps from distribution adjustments applied to prompts at different epochs, both before and after prompt tuning, on VTAB-1k Natural DTD using ViT-B/16. For the adding-based adjustment, we trained with a total of 59 prompts distributed uniformly across 12 Transformer blocks, resulting in one block containing 4 prompts while the others contained 5, then added a new prompt to the block with 4 prompts. For the repositioning-based adjustment, we trained with 60 prompts allocated uniformly, and then repositioned a single prompt. For fair comparison, we averaged the results over different configurations, including variations in which block had the missing prompt and which specific prompt was repositioned.
@@
Comparing the left and right parts of Fig. i, we find that merely adjusting the distribution without prompt tuning leads to negligible changes or even degrades performance. In contrast, implementing prompt tuning after adjusting the distribution leads to significant performance changes, with well-chosen adjustments leading to notable improvements (e.g., an increase of up to 2.2% achieved by repositioning just a single prompt). This brings us to:
Discovery 3. The effectiveness of distribution adjustments becomes apparent only after prompt tuning, suggesting that these adjustments can only be properly evaluated after tuning, thereby establishing a nested mechanism between ADO and VPT.
[…]
Discovery 2. Adjustments in prompt distribution are influenced by the updated prompts themselves, underscoring the necessity of an iterative mechanism that continuously refines the distribution over prompt updates.
@@
Additionally, to validate that the performance improvements from applying both distribution adjustments and prompt tuning stem primarily from the distribution adjustments rather than the tuning itself, we display the corresponding accuracy curve of prompt tuning in Fig. iii. Clearly, the prompted model has already converged around epoch 20, with only slight performance fluctuations thereafter (less than 0.3%). This indicates that the performance differences observed in Fig. i (typically exceeding 0.3%) cannot be attributed solely to prompt tuning alone, but rather to the adjustments in prompt distribution.
@@
Building upon the mechanisms described in Discoveries 2 and 3, a proper workflow for ADO and VPT should be structured as an iterative process, with VPT nested within ADO. Specifically, in each iteration, the process should first adjust the distribution to construct $\mathcal{D}^{\ast}$, followed by tuning the visual prompts to obtain $\mathbf{P}^{\ast}$. Based on $\mathcal{D}^{\ast}$ and $\mathbf{P}^{\ast}$, the distribution adjustments can be properly evaluated at the end. Fig. iii illustrates this co-design workflow for ADO and VPT. Formally, it can be expressed as a nested optimization problem, as follows:
$\mathcal{D}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathcal{D}}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\mathcal{L}(f_{\mathbf{P}^{\ast},\mathcal{D}}(\mathbf{x}),y)\big{]},$ (viii)
$\mathbf{P}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathbf{P}}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{\ast}}(\mathbf{x}),y)\big{]},$ (ix)
where the notations $\mathbf{P}^{\ast}$ and $\mathcal{D}^{\ast}$ correspond to those depicted in Fig. iii. This formulation is applicable to various distribution adjustment strategies, including adding, repositioning, and even pruning.
@@
Overall, the ADO problem is naturally formulated as a nested problem, grounded in its underlying mechanisms. Notably, since distribution adjustments can only be evaluated at the end, a prediction for Eq. viii is necessary for selecting appropriate adjustments. To this end, it is natural to frame ADO as a RL problem, where the reward is predicted to determine the next action, and the action’s effectiveness is evaluated afterward.
Appendix C Inferior Performance of Adding-based Adjustments
@@
Following the nested optimization framework described in Eqs. (viii) and (ix), we have also attempted to develop an adding-based strategy for ADO. Given that framing the ADO objective as a RL problem is a suitable choice, we also address the adding-based ADO by leveraging RL, with the components of the Markov decision process specified as follows:
@@
2) Action. Unlike prompt repositioning, which considers $L^{2}$ possible block arrangements, the adding-based ADO involves only $L$ potential adding operations. As a result, the adding-based ADO does not require simplification, and the action is straightforwardly represented as $a\in[L]$.
3) Reward. Similar to the PR strategy, the reward is formulated based on Eq. (viii) as $r\!=\!\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P}^{\prime},\mathcal{D}^{\prime}})$, where $\mathcal{D}^{\prime}$ denotes the updated prompt distribution after adding, and $\mathbf{P}^{\prime}$ represents the tuned prompts.
@@
Appendix D Relocating One Prompt for Each Iteration
Here, we explain why we restrict the relocation process to operate on only a single prompt for each iteration. Since we frame the allocation step as a RL problem, we need to evaluate the effectiveness of each individual allocation decision for relocated prompts. Relocating more than one prompt simultaneously would necessitate multiple reward evaluations, substantially increasing task complexity and computational overhead. Therefore, we designed to relocate only one prompt per iteration to maintain simplicity and efficiency in the allocation process.
Appendix E Derivation of the Taylor Expansion for the Idleness Score
Building upon the pruning objective in Eq. (3), the original idleness score is defined as:
@@ Appendix E Derivation of the Taylor Expansion for the Idleness Score
where $d_{k}=0$ indicates that the $k^{\text{th}}$ prompt is pruned. Specifically, if the idleness score $\mathcal{I}_{k}\!>\!0$, it indicates that the loss after pruning the prompt $\mathbf{p}_{k}$, i.e. $\mathcal{L}(f_{\mathbf{P},\mathcal{D}\mid d_{k}=0}(\mathbf{x}),y)$, is lower than the loss before pruning, i.e. $\mathcal{L}(f_{\mathbf{P},\mathcal{D}}(\mathbf{x}),y)$. This suggests that $\mathbf{p}_{k}$ has a negative impact on the current model $f_{\mathbf{P},\mathcal{D}}$ and can therefore be pruned for potential relocation.
Equivalently, this equation can be expressed as:
@@
where $\mathbf{p}_{k}=\mathbf{0}$ represents that the $k^{\text{th}}$ prompt is a zero vector.
@@
^$\begin{split}\widehat{\mathcal{I}}_{kj}&\approx\mathbf{g}^{T}\Big{(}\mathbf{P}-\mathbf{P}\!\mid\!p_{kj}\!=\!0\Big{)}\\ &\approx g_{kj}p_{kj},\end{split}$ (xiii)
^$\widehat{\mathcal{I}}_{k}\approx\sum_{j=1}^{d}g_{kj}p_{kj}\approx\mathbf{g}^{T}_{k}\mathbf{p}_{k}.$ (xiv)
Table ii: Specifications of the VTAB-1k datasets.
@@
Natural CIFAR-100 $100$ $800$ $200$ 10 0001000010\,00010 000
Caltech-101 $102$ 6 08460846\,0846 084
DTD $47$ 1 88018801\,8801 880
Oxford Flowers $102$ 6 14961496\,1496 149
Pets $73$ 66936693\,6693 669
SVHN $10$ 26 0322603226\,03226 032
Sun397 $397$ 21 7502175021\,75021 750
Specialized Patch Camelyon $2$ $800$ $200$ 32 7683276832\,76832 768
EuroSAT $10$ 5 40054005\,4005 400
RESISC45 $45$ 6 30063006\,3006 300
Diabetic Retinopathy $5$ 42 6704267042\,67042 670
Structured CLEVR-Count $8$ $800$ $200$ 15 0001500015\,00015 000
CLEVR-Distance $6$ 15 0001500015\,00015 000
DMLab $6$ 22 7352273522\,73522 735
KITTI-Distance $4$ $711$
dSprites-Location $16$ 73 7287372873\,72873 728
dSprites-Orientation $16$ 73 7287372873\,72873 728
smallNORB-Azimuth $18$ 12 1501215012\,15012 150
smallNORB-Elevation $9$ 12 1501215012\,15012 150
@@
CUB-200-2011* [51] $200$ 5 39453945\,3945 394 $600$ 5 79457945\,7945 794
NABirds* [50] $555$ 21 5362153621\,53621 536 2 39323932\,3932 393 6 08460846\,0846 084
Oxford Flowers [40] $021$ 02010201\,0201 020 1 02010201\,0201 020 6 14961496\,1496 149
Stanford Dogs* [6] $120$ 10 8001080010\,80010 800 1 20012001\,2001 200 8 58085808\,5808 580
Stanford Cars* [10] $196$ 7 32973297\,3297 329 $158$ 04180418\,0418 041
@@
Appendix F More Implementation Details
Augmentation. Apart from data normalization, we resize the input images to 224$\times$224 pixels for VTAB-1k and apply a randomly resize crop to 224$\times$224 pixels and horizontal flipping for FGVC, as outlined in [20, 13, 47].
Training Hyper-parameters. Specific to training hyper-parameters, we largely adopt the same settings as depicted in VPT [20]. Tab. i summarizes the hyper-parameter configurations comparing the experiments of VPT and our approach. Following [20], we conduct a grid search on the validation set of each task to determine the optimal learning rate and weight decay; the learning rate is set as $base\_lr\!\times\!b/\text{256}$, where $b$ denotes the batch size and $base\_lr$ is selected from the range specified in Tab. i. Notably, PRO-VPT does not require specific-designed large learning rates as in [20]. For all experiments conducted with our implementation, the results are averaged over three random seeds.
Appendix G More Experimental Results
@@
[…]
@@
Generalizability Study on Detection and Segmentation Tasks. We also conduct experiments on a broader range of downstream tasks, including object detection, instance segmentation, and semantic segmentation. Specifically, we evaluate object detection and instance segmentation on the COCO dataset, using Mask R-CNN with a Swin-T backbone pre-trained on ImageNet-1k. For semantic segmentation, we use the ADE20k dataset and adopt SETR-PUP with a ViT-B/16 backbone pre-trained on ImageNet-21k. As shown in Tab. vii, the results demonstrate that PRO-VPT consistently outperforms VPT across both detection and segmentation tasks, highlighting its superior generalizability.
Figure vii: Cost comparisons with different prompt numbers. We report the latency (ms/img) and the peak memory usage (GB). The increased training time associated with PRO-VPT is justified by the considerable enhancements it provides in both performance and robustness.
@@
Appendix H Visualization and Analysis
Figure viii: Visualization of the prompt distributions learned by PRO-VPT and the accuracy curves compared to VPT on the VTAB-1k datasets: Natural DTD, SVHN; Specialized Camelyon, Resisc45; and Structured DMLab, KITTI-Dist.

### v2

@@ Appendix A Detailed Explanations of Technical Concepts
where $\pi$ is the policy, $\theta$ is the policy parameters, and $k$ is the $k$-th step. It typically takes multiple steps of SGD to optimize the clipped surrogate function $L(\cdot,\cdot,\cdot,\cdot)$:
$L(s,a,\theta_{k},\theta)=$
$\;\min\big(\frac{\pi_{\theta}(a\!\mid\!s)}{\pi_{\theta_{k}}(a\!\mid\!s)}A^{\pi_{\theta_{k}}}(s,a),\text{clip}(\frac{\pi_{\theta}(a\!\mid\!s)}{\pi_{\theta_{k}}(a\!\mid\!s)},1\!-\!\epsilon,1\!+\!\epsilon)A^{\pi_{\theta_{k}}}(s,a)\big),$ (vii)
where $A^{\pi_{\theta_{k}}}(s,a)$ is the advantage estimator, and $\epsilon$ is the clip hyper-parameter. The clipping mechanism constrains the new policy to stay close to the old one, preventing excessively large policy updates that could degrade performance [35, 48].
Input: pre-trained model $f$, number of epochs $T$, number of prompts NN, learning rate $\eta$.
1 Initialize NN prompts $\mathbf{P}$ and distribute them to the model according to a uniform distribution $\mathcal{D}$.
2 for $t=0,\ldots,T-1$ do
3  Compute idleness scores $\{\widehat{\mathcal{I}}_{k}\}^{N}_{k=1}$ by Eq. (6).
4  if $\max\widehat{\mathcal{I}}_{k}\!>\!0$ then
5     Prune the negative prompt $\mathbf{p}_{k^{\!\ast\!}}$ with the maximum idleness score $\widehat{\mathcal{I}}_{k^{\!\ast\!}}$ to form $\mathcal{D}^{-}$.
6     Compute the current state $s$.
7     Compute the action $a^{\ast}\leftarrow\text{PPO}(s)$.
8     Allocate the idle prompt $\mathbf{p}_{k^{\!\ast\!}}$ to the $a^{\ast}$-th block to form $\mathcal{D}^{+}$.
9
10  end if
12  if $\max\widehat{\mathcal{I}}_{k}\!>\!0$ then
13     Compute the reward $\hat{r}$ by Eq. (8).
14     Update the policy networks within PPO based on $\hat{r}$.
15  end if
16
17 end for
Algorithm 1 PRO-VPT
Appendix B Pseudo-Code of PRO-VPT
We provide the detailed pseudo-code for PRO-VPT in Algo. 1. Initially, it generates $N$ prompts $\mathbf{P}$ and distributes them to the PVM according to a uniform distribution $\mathcal{D}$ (Line 1). In each epoch of PRO-VPT, it first calculates the estimated idleness scores $\{\widehat{\mathcal{I}}_{k}\}^{N}_{k=1}$ using Eq. (6) (Line 3). If $\max\widehat{\mathcal{I}}_{k}\!>\!0$, the PR process is triggered. This process begins by pruning the prompt $\mathbf{p}_{k^{\!\ast\!}}$ with the highest score $\widehat{\mathcal{I}}_{k^{\!\ast\!}}$, thereby constructing an intermediate distribution $\mathcal{D}^{-}$ (Line 5). The current state of the distribution $s$ is then computed, followed by determining the action $a^{\ast}$ based on PPO (Lines 6 and 7). Subsequently, the pruned prompt $\mathbf{p}_{k^{\!\ast\!}}$ is allocated to the $a^{\ast}$-th block, resulting in a relocated distribution $\mathcal{D}^{+}$ (Line 8). After completing the PR process, the prompts are optimized to $\mathbf{P}^{\prime}$ (Line 10). Additionally, if the PR process is activated during the epoch, it is necessary to calculate the estimated reward $\hat{r}$ using Eq. (8) and update the policy networks in PPO accordingly (Lines 12 and 13).
Appendix C Comprehensive Analysis of the Underlying Nature behind ADO
@@
Fig. i presents the complete results of Fig. 2. Specifically, we investigated the performance gaps from distribution adjustments applied to prompts at different epochs, both before and after prompt tuning, on VTAB-1k Natural DTD using ViT-B/16. For the adding-based adjustment, we trained with a total of 59 prompts distributed uniformly across 12 Transformer blocks, resulting in one block containing 4 prompts while the others contained 5, then added a new prompt to the block with 4 prompts. For the repositioning-based adjustment, we trained with 60 prompts allocated uniformly, and then repositioned a single prompt. For fair comparisons, we averaged the results over five runs with different configurations, including variations in which block had the missing prompt as well as which specific prompt was repositioned.
The following provides a summary of the analysis and key findings presented in the main paper:
@@
Finding 3. The effectiveness of distribution adjustments becomes apparent only after prompt tuning, suggesting that these adjustments can only be properly evaluated after tuning, thereby forming a nested relationship between ADO and VPT. Comparing the left and right parts of Fig. i, we find that adjusting the distribution without prompt tuning leads to negligible changes or even degrades performance. In contrast, implementing prompt tuning after adjusting the distribution leads to significant performance changes, with well-chosen adjustments leading to notable improvements (e.g., an increase of up to 2.2 pp achieved by repositioning just a single prompt). To this end, the proper workflow for ADO and VPT should be established as ‘distribution adjustment $\rightarrow$ prompt tuning $\rightarrow$ adjustment evaluation’, with prompt tuning nested within the distribution optimization process.
Finding 2. Adjustments in prompt distribution are influenced by the updated prompts themselves, underscoring the necessity of an iterative process that continuously refines the distribution over prompt updates. As demonstrated in the right portion of Fig. i, for prompts from epoch 25, we observe significant performance gains by adding a new prompt to the 10-th and 12-th blocks, as well as by repositioning one prompt to the 10-th and 11-th blocks. However, for prompts at epoch 75, the key improvements shifted to adding to the 6-th block and repositioning to the 6-th and 9-th blocks. To this end, the ADO-VPT co-design workflow should be established as an iterative process, which would be ‘distribution adjustment $\rightarrow$ prompt tuning $\rightarrow$ adjustment evaluation $\rightarrow$ new adjustment’.
@@
Additionally, to validate that the performance improvements from applying both distribution adjustments and prompt tuning mainly result from the distribution adjustments rather than the tuning itself, we present the respective prompt tuning accuracy curve in Fig. iii. Clearly, the prompted model has already converged around epoch 20, with only slight performance fluctuations thereafter (less than 0.3 pp). This indicates that the performance differences observed in Fig. i (typically exceeding 0.3 pp) cannot be attributed solely to prompt tuning alone, but rather to the adjustments for the prompt distribution.
Figure $i$: Convergence curve across different prompt-tuning epochs, which corresponds to Fig. i.
@@
Building upon the underlying nature described in Findings 2 and 3, a proper workflow for ADO and VPT should be structured as an iterative nested process. Specifically, in each iteration, the process first adjusts the distribution to construct $\mathcal{D}^{\ast}$, then tunes the visual prompts to obtain $\mathbf{P}^{\ast}$. Based on $\mathcal{D}^{\ast}$ and $\mathbf{P}^{\ast}$, the effectiveness of the distribution adjustment is evaluated at the end of each iteration, marking the completion of the current cycle and the beginning of the next. Fig. iii illustrates this co-design workflow for ADO and VPT. Formally, it can be expressed as a nested optimization problem as follows:
$\mathcal{D}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathcal{D}}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\mathcal{L}(f_{\mathbf{P}^{\ast},\mathcal{D}}(\mathbf{x}),y)\big],$ (viii)
$\mathbf{P}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathbf{P}}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{\ast}}(\mathbf{x}),y)\big],$ (ix)
where the notations $\mathbf{P}^{\ast}$ and $\mathcal{D}^{\ast}$ correspond to those depicted in Fig. iii.
@@
Overall, the ADO problem is naturally formulated as a nested problem, grounded in its underlying nature. This formulation is applicable to various distribution adjustment strategies (e.g., adding, repositioning, or even pruning) as well as different discrete optimization methods (e.g., evolutionary algorithms and reinforcement learning). Notably, since distribution adjustments can only be evaluated at the end, a prediction for Eq. viii is necessary for selecting an appropriate adjustment. To this end, it is natural to frame ADO as a RL problem, where the reward is predicted to determine the next action, and the effectiveness of that action is evaluated afterward.
Appendix D Inferior Performance of Adding-based Adjustments
@@
Following the nested optimization framework described in Eqs. (viii) and (ix), we have also attempted to develop an adding-based strategy for ADO. Given that framing the ADO objective as a RL problem is a suitable choice (refer to § 3.3 and § C), we also address the adding-based ADO by leveraging RL, with the components of the Markov decision process specified as follows:
@@
2) Action. Unlike prompt repositioning, which considers $L^{2}$ possible block arrangements, the adding-based ADO involves only $L$ potential adding operations. As a result, the adding-based ADO does not require decomposition for RL, and the action is straightforwardly represented as $a\in[L]$.
3) Reward. Similar to the PR strategy, the reward is formulated based on Eq. (viii) as $r\!=\!\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P}^{\prime},\mathcal{D}^{\prime}})$, where $\mathcal{D}^{\prime}$ denotes the updated prompt distribution after adding and $\mathbf{P}^{\prime}$ represents the tuned prompts.
@@
Appendix E Relocating One Prompt for Each Iteration
Here, we explain why we restrict the relocation process to operate on only a single prompt for each iteration. Since we frame the allocation step as a RL problem, we need to evaluate the effectiveness of each allocation decision for relocated prompts. Relocating more than one prompt simultaneously would necessitate multiple reward evaluations, substantially increasing task complexity and computational overhead. Therefore, we designed to relocate only one prompt per iteration to maintain simplicity and efficiency in the allocation process.
Appendix F Derivation of the Taylor Expansion for the Idleness Score
Based on the pruning objective in Eq. (3), the original idleness score is defined as:
@@ Appendix F Derivation of the Taylor Expansion for the Idleness Score
where $d_{k}=0$ indicates that the $k$-th prompt is pruned. Equivalently, this equation can be expressed as:
@@
where $\mathbf{p}_{k}=\mathbf{0}$ represents that the $k$-th prompt is a zero vector.
@@
$\begin{split}\widehat{\mathcal{I}}_{kj}&\approx\mathbf{g}^{T}\Big(\mathbf{P}-\mathbf{P}\!\mid\!p_{kj}\!=\!0\Big)\\ &\approx g_{kj}p_{kj},\end{split}$ (xiii)
$\widehat{\mathcal{I}}_{k}\approx\sum_{j=1}^{d}g_{kj}p_{kj}\approx\mathbf{g}^{T}_{k}\mathbf{p}_{k}.$ (xiv)
Appendix G More Implementation Details
Data Augmentation. Apart from data normalization, we resize the input images to $\times224$ pixels for VTAB-1k and apply a randomly resize crop to $\times224$ pixels and horizontal flipping for FGVC, as outlined in [22, 13, 51, 59].
@@ Appendix G More Implementation Details
Training Hyper-Parameters. Specific to training hyper-parameters, we largely adopt the same settings as depicted in VPT [22]. Tab. i summarizes the hyper-parameter configurations comparing the experiments of VPT and our approach. Following [22, 13], we conduct a grid search on the validation set of each task to determine the optimal learning rate and weight decay; the learning rate is set as $base\_lr\!\times\!b/\text{256}$, where $b$ denotes the batch size and base_lrbase\_lr is selected from the range specified in Tab. i. Notably, PRO-VPT does not require specific-designed large learning rates as in [22]. For all experiments conducted with our implementation, the results are averaged over three random seeds.
Table $i$: Specifications of the VTAB-1k datasets.
@@
Natural CIFAR-100 $100$ $800$ $200$ 10 00010\,000
Caltech-101 $102$ 6 0846\,084
DTD $47$ 1 8801\,880
Oxford Flowers $102$ 6 1496\,149
Pets $37$ 3 6693\,669
SVHN $10$ 26 03226\,032
Sun397 $397$ 21 75021\,750
Specialized Patch Camelyon 22 $800$ $200$ 32 76832\,768
EuroSAT $10$ 5 4005\,400
RESISC45 $45$ 6 3006\,300
Diabetic Retinopathy 55 42 67042\,670
Structured CLEVR-Count 88 $800$ $200$ 15 00015\,000
CLEVR-Distance 66 15 00015\,000
DMLab 66 22 73522\,735
KITTI-Distance 44 $711$
dSprites-Location $16$ 73 72873\,728
dSprites-Orientation $16$ 73 72873\,728
smallNORB-Azimuth $18$ 12 15012\,150
smallNORB-Elevation 99 12 15012\,150
Appendix H More Experimental Results
@@
CUB-200-2011* [56] $200$ 5 3945\,394 $600$ 5 7945\,794
NABirds* [55] $555$ 21 53621\,536 2 3932\,393 6 0846\,084
Oxford Flowers [42] $102$ 1 0201\,020 1 0201\,020 6 1496\,149
Stanford Dogs* [6] $120$ 10 80010\,800 1 2001\,200 8 5808\,580
Stanford Cars* [10] $196$ 7 3297\,329 $815$ 8 0418\,041
@@
APb $^{b}_{\text{50}} AP$$^{b}_{\text{75}}$ APm $^{m}_{\text{50}} AP$$^{m}_{\text{75}}$ mIoU-$S$ mIoU-MS
@@
Generalizability Study on Detection and Segmentation Tasks. We also conduct experiments on a broader range of downstream tasks, including object detection, instance segmentation, and semantic segmentation. Specifically, we evaluate object detection and instance segmentation performance on the COCO dataset [36], using Mask R-CNN [17] with a Swin-T backbone pre-trained on ImageNet-1k. For the semantic segmentation task, we use the ADE20k dataset [70] and adopt SETR-PUP [69] with a ViT-B/16 backbone pre-trained on ImageNet-21k. As shown in Tab. vii, the results demonstrate that PRO-VPT consistently outperforms VPT across both detection and segmentation tasks, highlighting its superior generalizability.
Figure vii: Cost comparisons with different prompt numbers. We report the latency (ms/img) and the GPU memory usage (GB). The increased training time associated with PRO-VPT is justified by the considerable enhancements in both performance and robustness.
@@
Appendix I Visualization and Analysis